2)PQR is a straight line PQ=4PR P(-4,-1) Q(6,11) find the co-ordinate of R?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39618)

Fix the description. PQ is less than PR because Q is between P and R. Either PQR is the line using Q between P and R; or PQ is some fraction of PR. Both cannot be true at the same time.

Answer by stanbon(75887)

PQR is a straight line PQ=4PR P(-4,-1) Q(6,11) find the co-ordinate of R
--------------
length of PQ = sqrt[(11--1)^2+(6--4)^2] = sqrt[144+100)
= sqrt(244) = 2sqrt(61)
-----
PR = (1/4)PQ = (1/2)sqrt(61)
----
Comment: since PQ is larger than PR, R must be between P and Q
----
Ratio of PQ to PR = 2sqrt(61)/(1/2)sqrt(61) = 4:1
-----
So R is 1/4 of the way from P to Q
---
Let the coordinates of R be (x,y)
x = (1/4)(6--4) = (1/4)(10) = 2.5
y = (1/4)(11--1) = (1/4)(12) = 3
---
Ans:: R = (2.5,3)
-------------
Cheers,
Stan H.
-------------